# View-Invariant Pixelwise Anomaly Detection in Multi-object Scenes with Adaptive View Synthesis

### v1

@@ Abstract
The inspection and monitoring of infrastructure assets typically requires identifying visual anomalies in scenes periodically photographed over time. Images collected manually or with robots such as unmanned aerial vehicles from the same scene at different instances in time are typically not perfectly aligned. Supervised segmentation methods can be applied to identify known problems, but unsupervised anomaly detection approaches are required when unknown anomalies occur. Current unsupervised pixel-level anomaly detection methods have mainly been developed for industrial settings where the camera position is known and constant. However, we find that these methods fail to generalize to the case when images are not perfectly aligned. We term the problem of unsupervised anomaly detection between two such imperfectly aligned sets of images as Scene Anomaly Detection (Scene AD). We present a novel network termed OmniAD to address the Scene AD problem posed. Specifically, we refine the anomaly detection method reverse distillation to achieve a 40% increase in pixel-level anomaly detection performance. The network’s performance is further demonstrated to improve with two new data augmentation strategies proposed that leverage novel view synthesis and camera localization to improve generalization. We validate our approach with qualitative and quantitative results on a new dataset, ToyCity, the first Scene AD dataset with multiple objects, as well as on the established single object-centric dataset, MAD. https://drags99.github.io/OmniAD/
Index Terms:
@@
Anomaly Detection, Unsupervised Learning, Camera Localization
I Introduction
@@
Figure 1: In Scene AD, we observe variations in camera positions between non-anomalous views (represented in green), and query views which may have anomalies in them (represented in red). In contrast, traditional AD methods typically feature aligned views between the non-anomalous views and anomalous views.
The inspection and monitoring of infrastructure assets typically requires identifying visual anomalies in scenes periodically photographed over time [1, 2, 3, 4, 5]. Images collected manually or with robots such as unmanned aerial vehicles from the same scene at different instances in time are typically not perfectly aligned. For instance, the United States Navy places significant emphasis on the monitoring and detection of anomalies during inspections of their ships, which is critical for maintaining security and operational readiness and safety [6, 7]. Similarly, bridge inspectors can leverage this technology to identify parts of the bridge that are changing or behaving abnormally. This information is vital for predicting future performance and for maintenance decision making. In each of these examples, new image sets are acquired at every inspection but typically unaligned with previous inspections.
We define the challenge of detecting anomalous regions between a reference non-anomalous set of images and a recapture of the scene with unaligned views as Scene Anomaly Detection (Scene AD). The recaptured scene will be referred to as the query scene and the corresponding set of images as the query set. Importantly, the query scene may or may not contain any anomalies. Additionally, what may constitute as an anomalous region is not available in most scenarios, thus providing supervision for supervised methods is usually unfeasible. Therefore, Scene AD requires methods that do not rely on any supervision for what may constitute as an anomalous region.
Prior unsupervised anomaly detection methods have not been designed to handle Scene AD and are typically developed for controlled industrial settings with constant camera positions. As a result, we find that state-of-the art anomaly detection methods perform poorly on the Scene AD task. Apart from the architectures themselves, we find that there are two main challenges that need to be overcome. First, there is no benchmark multi-object dataset available for Scene AD. As described earlier, existing AD benchmarks typically focus on a single view. Zhou et al. proposed a multi-pose AD [8] dataset, however this problem setting focuses on multiple poses of a single object and not that of an entire scene. Second, any solution to Scene AD must thus incorporate a mechanism for view-invariance.
Due to the limited generalizability of prior unsupervised anomaly detection (AD) methods in Scene AD, we propose OmniAD, a refinement of Reverse Distillation (RD) [9]. To detect anomalous region locations using RD, deviations in a student decoder must occur with respect to a teacher encoder. We postulate that expanding the effective receptive field (ERF) [10] of the student decoder will increase the deviations produced by the student decoder, which in turn will increase the sensitivity to anomalous regions. To expand the ERF, we incorporate student attention modules into the student decoder. We show quantitatively that the expansion of the ERF will, in turn, increase the performance of RD.
However, to achieve improved view-invariance, we introduce two new data augmentation methods. Building on the intuition that it is much easier to compare perfectly aligned pictures, rather than comparing misaligned ones, our data augmentation methods utilize novel view synthesis (NVS) techniques. Specifically, we employ Neural Radiance Fields (NeRF) [11], combined with a camera localization strategy, hierarchical localization (hloc) [12, 13], to generate novel views of the query scene that closely align with the non-anomalous views. We hypothesize that providing these generated aligned novel views will allow AD methods to improve generalizability. Although supervised methods for anomaly detection cannot leverage this data augmentation due to the lack of ground truth for anomalies, unsupervised methods can effectively use this approach for improved performance.
To properly evaluate and benchmark the effectiveness of these methods, a suitable dataset is required. Recognizing the lack of a comprehensive benchmark for Scene AD in real scenes with multiple objects, we introduce the ToyCity dataset. ToyCity is tailored to provide a robust real benchmark featuring multiple camera viewpoints and a range of anomalies across different setups. ToyCity’s unique structure presents a new and significant challenge to state-of-the-art AD models. We evaluate several AD methods, including OmniAD, using ToyCity as well as an object focused dataset, the Multi-pose Anomaly Detection dataset (MAD) [8].
To summarize our main contributions are:
  • Introduction of the novel computer vision task of Scene AD involving detecting anomalous regions between a reference non-anomalous set of images and a recapture of the scene with unaligned views. Results are presented to demonstrate the inability of existing anomaly detection methods to generalize effectively to the Scene AD task.
  • The ToyCity dataset specifically designed for Scene AD, addressing the need for suitable benchmarks for multi-object and multi-view problems in this domain.
  • Development of a new anomaly detection architecture termed OmniAD that better generalizes to Scene AD. Specifically, we introduce a novel Student Attention Module to expand the effective receptive field (ERF) of the reverse distillation (RD) architecture and demonstrate state-of-the art performance on the ToyCity and MAD datasets [8].
  • Two novel data augmentation techniques leveraging Neural Radiance Fields and hierarchical localization, resulting in a 40% increase in performance for pixel-wise localization of anomalous regions in OmniAD.
This paper is organized as follows: II provides a review of the relevant literature in the area of anomaly detection, neural radiance fields, and hierarchical localization. III describes the proposed methodology for data augmentation as well as OmniAD. IV discusses the structure of the ToyCity dataset, developed to evaluate AD methods for scene AD. V describes the experimental setup, evaluation metrics, and both a quantitative and qualitative analysis of the experimental results. Finally, VI is a conclusion to the paper, summarizing the main findings, followed by references.
II Related Work
@@
III Method
An overview of our AD methodology for Scene AD can be seen in Figure2. Scene AD consists of non-anomalous views and query views. We generate an SfM model using non-anomalous views to acquire camera poses. Leveraging these non-anomalous camera poses, we can implement two different novel view selection strategies to produce images for data augmentation as shown in Figure3. For the first strategy, we interpolate between the non-anomalous camera poses. The second strategy for generating useful camera views for training is to utilize hloc to acquire the camera poses of query views. We then use NVS with the new camera poses to augment our non-anomalous views for training our proposed method OmniAD. The following subsections describes each of these steps in more detail.
@@ III Method
III-A Novel View Selection Strategy
@@
III-A1 Interpolated Non-anomalous Views
@@
Given a set of non-anomalous images an associated camera poses $T$ can be calculated using SfM. The calculated $T$ can be interconnected to form a trajectory. A simple approach for this interconnection is a greedy algorithm that selects a starting camera position and sequentially moves to the nearest neighboring camera position. Specifically, for our experiments, we interpolate 12 intermediate camera poses between consecutive original poses in $T$, thereby enhancing the dataset with interpolated views that increase the robustness of our AD model to various camera angles. An example of the camera interpolation for ToyCity can be seen in Figure4.
III-A2 Query-Aligned Non-anomalous Views
@@
III-B Novel View Synthesis
@@
Once the new camera poses are selected, we construct our augmented datasets using NVS, as shown in Figure3. Example camera interpolated images can be seen in Figure3, and example images using QANV can be seen in Figure3
III-C OmniAD
We develop OmniAD to leverage our proposed augmentation techniques and surpass the performance of state-of-the-art AD methods when applied to Scene AD. OmniAD significantly enhances the RD framework through two significant updates, (i) integrating a ResNeXt backbone in the teacher encoders[33] (ii) and adding novel student attention modules, $A^{i}$. We found a combination of the ResNext with the student attention module to offer a performance increase of up to 40%, while the training and inference time cost for OmniAD is similar to RD.
@@ III-C OmniAD
Student attention modules $A^{i}$ are introduced after each of the Student Decoder networks $D^{i}$ of the RD architecture. These modules are designed to bolster OmniAD’s sensitivity to anomalous samples. An illustrative overview of the OmniAD architecture is presented in Figure5. $A^{i}$ is based on the U-Net architecture and consists of alternating layers of 3x3 convolutions with self attention layers and a bottleneck layer consisting of several self attention layers as shown in Figure6. This architecture was designed to expand the Effective Receptive Field (ERF) of the Student Decoder which in turn would increase the variations in the output features in response to an anomaly. We validate this hypothesis and visualize the larger ERF in Figure7. This means there is a larger perturbation of features when an anomalous region is introduced into the model, making it more sensitive and responsive to such anomalies. This increased sensitivity is attributed to the improved performance of OmniAD. We conducted experiments to validate this hypothesis and demonstrate the efficacy of $A^{i}$.
@@
IV ToyCity Dataset
The ToyCity dataset addresses the need for a benchmark dataset consisting of real images of multiple objects for Scene AD. Unlike MAD, ToyCity consists of (i) substantially more real images and (ii) is not centered around a single object. ToyCity features scenes of a toy city with buildings, signs, roads, trees, and miscellaneous items. The dataset contains images from various angles and heights in four distinct scene arrangements, as illustrated in Figure8. Each setup also contains different anomalies, as shown in Figure8. Additoinally, each anomaly image has at most one anomaly. The dataset includes the manually annotated ground truth for anomaly segmentation. Table I provides an overview of the number of images released for ToyCity. The dataset’s v1 is the base variant of all four scenes without any augmentations, v2 includes Quer-Aligned Novel Views (QANV) images rendered using Nerfacto, v3 incorporates augmentations using Interpolated Non-Anomalous Views (INV), and v4 combines the data of both v2 and v3, these variations are presented in Table II. ToyCity thus offers a new benchmarking challenge to current state-of-the-art anomaly detection models.
@@ IV ToyCity Dataset
TABLE I: Number of non-anomalous and anomalous images in each of the ToyCity dataset scenes, and the number of NVS images generated from each of the proposed methods QANV and IBV
@@
V Experiments
We evaluate several AD methods on ToyCity and MAD to test the localization ability of each method. We could only utilize a subset of the MAD dataset that consisted of real images compatible with COLMAP for Structure from Motion (SfM). Each AD method is evaluated quantitatively using the pixel-level $F_{1}$ score and AUROC. An ablation study follows to show the significance of the student attention modules on RD. Finally a qualitative comparison between OmniAD and RD for anomaly segmentation is presented.
V-A Metrics
To comprehensively assess the anomaly localization performance of various AD models, we employ the pixel-wise $F_{1}$ Score as described by Equation 1. The $F_{1}$ score excels in handling class imbalances by merging precision and recall into one comprehensive metric. This balance is critical for datasets such as ToyCity and MAD, where anomaly regions typically constitute a small fraction of the total image area. Examples of the relatively small anomaly regions within ToyCity can be seen in Figure8, underscoring the necessity for such a balanced evaluation metric. The imbalance of anomaly regions renders other metrics like accuracy or the area under the receiver operating characteristic curve (AUROC) as misleading.

### v2

# OmniAD: View-Invariant Unsupervised Scene Pixelwise Anomaly Detection with Adaptive View Synthesis

@@ Abstract
Visual anomaly detection in the built environment can be a valuable tool for applications such as infrastructure assessment, construction monitoring, security surveillance, and urban planning. Anomaly detection approaches are typically unsupervised and work by detecting deviations from an expected state where no assumptions are made exact type of deviation. Unsupervised pixel-level anomaly detection methods have been developed to successfully recognize and segment anomalies; however, existing techniques are designed for industrial settings with a fixed camera position. In the built environment, images may be periodically collected by a moving camera manually or by aerial or ground vehicles and utilized for decision making. The camera pose between successive collections may vary widely voiding a fundamental assumption in existing anomaly detection approaches. To address this gap, we introduce the problem of Scene Anomaly Detection (Scene AD), where the goal is to detect anomalies from two sets of images: one set without anomalies and one set that may or may not contain anomalies. No labeled semantic segmentation data are provided for training. We propose a novel network, OmniAD, to tackle Scene AD by refining the reverse distillation anomaly detection method, leading to a 40% improvement in pixel-level anomaly detection. Additionally, we introduce two new data augmentation strategies that leverage novel view synthesis and camera localization to enhance generalization. We evaluate our approach both qualitatively and quantitatively on a new dataset, ToyCity the first Scene AD dataset featuring multiple objects as well as on the established single object centric dataset, MAD. Our method demonstrates marked improvement over baseline approaches, paving the way for robust anomaly detection in scenes with real-world camera pose variations commonly observed in the built environment. Project Page: https://drags99.github.io/OmniAD/
keywords:
@@
Anomaly Detection , Unsupervised Learning , Camera Localization , Neural Radiance Fields
[…]
[1] organization=Department of Electrical and Computer Engineering, University of Houston, addressline=4226 Martin Luther King Blvd., city=Houston, state=Texas, postcode=77204, country=USA
[…]
[2] organization=Department of Civil and Environmental Engineering, University of Houston, addressline=4226 Martin Luther King Blvd., city=Houston, state=Texas, postcode=77204, country=USA
1 Introduction
@@
Figure 1: In Scene AD, we observe variations in camera positions between non-anomalous views (green), and query views which may contain anomalies (red). In contrast, traditional AD methods typically feature aligned views between the non-anomalous and anomalous images.
For many applications in the built environment —ranging from infrastructure assessment and construction monitoring to security surveillance and urban planning—detecting visual anomalies in scenes captured over time is a critical task  [1, 2, 3, 4, 5]. These anomalies may arise from structural changes, newly introduced objects, unexpected modifications in appearance, or more generally, anything that deviates from an established non-anomalous reference. However, images acquired at different instances, whether manually or autonomously, typically suffer from misalignment due to variations in camera pose. For example, bridge inspectors may assess changes to a structure following a disaster, self-driving car systems might detect modifications to traffic signs, and the U.S. Military places strong emphasis on monitoring infrastructure for anomalies critical to operational readiness [6, 7, 8, 9, 10, 11, 12]. In all such cases, each new set of images is unaligned with previous captures, complicating anomaly identification.
We define the task of Scene Anomaly Detection (Scene AD) as detecting anomalous regions in a query set of images, potentially containing no anomalies at all, against a non-anomalous reference set. Labeled examples of what qualifies as an anomaly are unavailable, rendering supervised segmentation methods impractical. Additionally anomalies are often too infrequent or unpredictable to compile sufficient labeled training data for every potential anomaly type, rendering fully supervised models infeasible. As a result, unsupervised anomaly detection (AD) approaches are a feasible alternative. However, these AD methods were originally crafted for controlled industrial scenarios which commonly assume fixed camera views and struggle with the viewpoint variations of real-world Scene AD.
Two main challenges emerge. First, there is no suitable multi-object, multi-view benchmark for Scene AD: most existing AD datasets capture a single view or focus on a single object [13]. Second, any robust Scene AD solution must tolerate significant camera misalignment and still detect newly introduced or changed regions.
To address these gaps, we propose OmniAD, a refinement of the Reverse Distillation (RD) framework [14]. RD detects anomalies by comparing a student decoder’s reconstruction with a teacher encoder’s embedding. We hypothesize that broadening the student decoder’s effective receptive field (ERF) [15] will improve sensitivity to anomalous regions. Accordingly, we incorporate student attention modules to expand the ERF and enhance pixel-level anomaly detection.
Moreover, to manage viewpoint differences in real scenes, we introduce two data augmentation strategies leveraging Neural Radiance Fields (NeRF) [16] and hierarchical localization (hloc) [17, 18]. These techniques generate novel, more closely aligned views of the query scene, aiding unsupervised methods—such as OmniAD—in detecting anomalies without labeled examples.
Since no comprehensive benchmark exists for this multi-object, multi-view setting, we introduce the ToyCity dataset. ToyCity consists of scenes with multiple objects, varied camera viewpoints, and different anomalies. We evaluate OmniAD and several baseline methods on ToyCity as well as on the single-object-centric Multi-pose Anomaly Detection dataset (MAD) [13], demonstrating notable performance gains.
Our main contributions are:
  • 1.
    Scene AD Task: Formal introduction of Scene Anomaly Detection, highlighting the challenge of detecting anomalies in unaligned multi-view imagery.
  • 2.
    ToyCity Dataset: A new multi-object, multi-view dataset designed for benchmarking Scene AD methods in realistic settings.
  • 3.
    OmniAD Architecture: A refined Reverse Distillation approach with a proposed Student Attention Module to expand the ERF, achieving state-of-the-art performance on ToyCity and MAD [13] and validated with ablation studies.
  • 4.
    NeRF-based Data Augmentation Methods: Two novel techniques using NeRF and hierarchical localization for improved view-invariance, yielding a 40% increase in anomaly localization accuracy.
The rest of the paper is organized as follows: Section 2 reviews related work on anomaly detection, neural radiance fields, and camera localization. Section 3 introduces the proposed OmniAD architecture and novel data augmentation strategies. Section 4 presents the ToyCity dataset. Section 5 details the experimental setup, metrics, and results. Section 6 discusses the limitations of our dataset and method. Finally, Section 7 concludes the paper.
2 Related Work
@@
3 Method
An overview of our AD methodology for Scene AD can be seen in Figure  2. Scene AD consists of non-anomalous views and query views. We generate an SfM model using non-anomalous views to acquire camera poses. Leveraging these non-anomalous camera poses, we can implement two different novel view selection strategies to produce images for data augmentation as shown in Figure  3. For the first strategy, we interpolate between the non-anomalous camera poses. The second strategy for generating useful camera views for training is to utilize hloc to acquire the camera poses of query views. We then use NVS with the new camera poses to augment our non-anomalous views for training our proposed method OmniAD. The following subsections describes each of these steps in more detail.
@@ 3 Method
3.1 Novel View Selection Strategy
@@
3.1.1 Interpolated Non-anomalous Views
@@
Given a set of non-anomalous images an associated camera poses $T$ can be calculated using SfM. The calculated $T$ can be interconnected to form a trajectory. A simple approach for this interconnection is a greedy algorithm that selects a starting camera position and sequentially moves to the nearest neighboring camera position. Specifically, for our experiments, we interpolate 12 intermediate camera poses between consecutive original poses in $T$, thereby enhancing the dataset with interpolated views that increase the robustness of our AD model to various camera angles. An example of the camera interpolation for ToyCity can be seen in Figure  4.
3.1.2 Query-Aligned Non-anomalous Views
@@
3.2 Novel View Synthesis
@@
Once the new camera poses are selected, we construct our augmented datasets using NVS, as shown in Figure  3. Example camera interpolated images can be seen in Figure  3, and example images using QANV can be seen in Figure  3
3.3 OmniAD
Our goal in Scene AD is to handle significant camera misalignment while detecting anomalies at a pixel level. Although Reverse Distillation (RD) [14] provides a robust teacher student mechanism for fixed view scenarios, our preliminary tests revealed its limitations in multi-view scenarios.
Backbone Exploration. We first explored different backbones to replace RD’s default encoder, focusing on architectures that could accommodate the diverse viewpoints. Many alternatives struggled to converge or offered minimal gains, but ResNeXt [38] consistently yielded better multi-scale representations and improved performance compared to simpler encoders. In principle, any sufficiently expressive architecture including fully attention-based backbones could be substituted for the encoder. However, given the ever-growing variety of such networks, we restricted our exploration to widely adopted backbones and found that ResNeXt provided the most consistent gains for our multi-view setting.
Receptive Field Constraints. Further analysis indicated that RD’s student decoder had a relatively small effective receptive field (ERF) [15], causing it to overlook anomalies. This shortcoming proved critical in multi-view settings, where anomalies could appear in various positions or at different scales.
Rationale of ERF Expansion for Anomaly Sensitivity. Let $\mathcal{I}$ denote the image domain (e.g., $H\times W$ pixels), and let $x\in\mathbb{R}^{C\times H\times W}$ be an input image. Suppose a teacher network $T$ extracts multi-scale features
$T(x)\;=\;\bigl{(}T^{1}(x),T^{2}(x),\dots,T^{L}(x)\bigr{)},$
where $T^{l}(x)\in\mathbb{R}^{C_{l}\times H_{l}\times W_{l}}$ is the $l$-th feature map. A student network $S$ decodes these teacher features to produce a reconstruction $\hat{x}=(S\circ T)(x)$. In Reverse Distillation [14], any mismatch between $x$ and $\hat{x}$ can indicate an anomaly, since $S$ is trained on non-anomalous (normal) data alone.
Effective Receptive Field (ERF). For each output location $(i,j)$ in $\hat{x}$, define its effective receptive field as
$\mathrm{ERF}_{ij}(S\circ T)\;=\;\bigl{\{}\,(u,v)\in\mathcal{I}\;\big{|}\;\tfrac{\partial\hat{x}_{ij}}{\partial x_{uv}}\neq 0\bigr{\}}.$
In standard RD, convolutions and limited skip connections yield a relatively localized ERF. If a true anomaly lies outside $\mathrm{ERF}_{ij}$ for many $(i,j)$, it may not sufficiently alter $\hat{x}$, causing the anomaly to go undetected.
Impact of ERF Expansion. Let $A\subset\mathcal{I}$ denote the anomaly region, i.e., pixels in $x$ that deviate from normal. By adding student attention modules (Fig. 6), the new student $S^{\prime}$ broadens $\mathrm{ERF}_{ij}(S^{\prime}\circ T)$:
$\bigl{|}\mathrm{ERF}_{ij}(S^{\prime}\circ T)\bigr{|}\;\;\geq\;\;\bigl{|}\mathrm{ERF}_{ij}(S\circ T)\bigr{|}.$
Consequently, for any anomalous pixel $(u,v)\in A$, there are more $(i,j)$ in the reconstructed output $\hat{x}$ for which $(u,v)\in\mathrm{ERF}_{ij}(S^{\prime}\circ T)$. In other words, the anomaly influences a broader swath of the reconstruction, resulting in larger or more easily detectable errors.
Reverse Distillation Anomaly Score. As in conventional RD, we define an anomaly score per pixel $(i,j)$ by measuring discrepancy between $\hat{x}_{ij}$ and $x_{ij}$. When the ERF is small, an anomaly’s influence might be confined or even “explained away” by the network averaging out the region, yielding low detection sensitivity. With an expanded ERF, the anomaly is more likely to cause a significant mismatch for at least one output location $(i,j)$, thereby increasing detection rates.
Attention Modules for the Student Decoder. To address the ERF limitation, we introduced additional attention layers, denoted $A^{i}$, after each decoder block. These modules expand the ERF by allowing the network to aggregate information from spatially distant regions, thus capturing subtle or pose-dependent anomalies. Figure 6 shows the basic layout of each attention module, incorporating interleaved 3x3 convolutions and self-attention blocks for both local and global feature integration.
@@ 3.3 OmniAD
4 ToyCity Dataset
The ToyCity dataset addresses the need for a benchmark dataset consisting of real images of multiple objects for Scene AD. Unlike MAD, ToyCity consists of (i) substantially more real images and (ii) is not centered around a single object. ToyCity features scenes of a toy city with buildings, signs, roads, trees, and miscellaneous items. The dataset contains images from various angles and heights in four distinct scene arrangements, as illustrated in Figure  8. Each setup also contains different anomalies, as shown in Figure  8. Additionally, each anomaly image has at most one anomaly but each setup has varying anomalies that were added and removed. The dataset includes the manually annotated ground truth for anomaly segmentation. Table 1 provides an overview of the number of images released for ToyCity. The dataset’s v1 is the base variant of all four scenes without any augmentations, v2 includes Quer-Aligned Novel Views (QANV) images rendered using Nerfacto, v3 incorporates augmentations using Interpolated Non-Anomalous Views (INV), and v4 combines the data of both v2 and v3, these variations are presented in Table 2. ToyCity thus offers a new benchmarking challenge to current state-of-the-art anomaly detection models.
@@ 4 ToyCity Dataset
Table 1: Number of non-anomalous and anomalous images in each of the ToyCity dataset scenes, and the number of NVS images generated from each of the proposed methods QANV and IBV
@@
5 Experiments
We evaluate OmniAD’s performance, its ablated variants, and several AD methods on ToyCity and MAD to test the localization ability of each method. We could only utilize a subset of the MAD dataset that consisted of real images compatible with COLMAP for Structure from Motion (SfM). Each AD method is evaluated quantitatively using the pixel-level $F_{1}$ score and AUROC. An ablation study follows to show the significance of the student attention modules on RD. Finally a qualitative comparison between OmniAD and RD for anomaly segmentation is presented.
5.1 Metrics
To comprehensively assess the anomaly localization performance of various AD models, we employ the pixel-wise $F_{1}$ Score as described by Equation 1. The $F_{1}$ score excels in handling class imbalances by merging precision and recall into one comprehensive metric. This balance is critical for datasets such as ToyCity and MAD, where anomaly regions typically constitute a small fraction of the total image area. Examples of the relatively small anomaly regions within ToyCity can be seen in Figure  8, underscoring the necessity for such a balanced evaluation metric. The imbalance of anomaly regions renders other metrics like accuracy or the area under the receiver operating characteristic curve (AUROC) as misleading.